In manufacturing its product, City Manufacturing’s recipe requires 3 ounces of Material X at a standard cost of $0.20 per ounce. During January 2010, City bought and used 2,000 ounces of Material X at a cost of $0.18 per ounce and made 700 units of product. The company has estimated that it will produce 750 units of product per month.What is City's material price variance?
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (0.20 - 0.18) * 2,000

= 0.02 * 2000

= 40 FAVOURABLE